# The commutativity of cofinal types

### v1

@@ 1.2. The cofinal type of sums and products
$({\dagger})\ \ \text{whenever }Y\subseteq X,\text{ and }C\in U\times\prod_{\alpha\in X}V_{\alpha},\ f_{Y}(\pi_{X,Y}(C))\geq_{\mathbb{P}}f_{X}(C).$
@@
Suppose that $\mathbb{P}$ is a complete order. Then $\mathbb{P}$ is uniformly bellow $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
@@
[…]
@@
Suppose that $\mathbb{P}$ is complete and for each $X\in U$, $\mathcal{X}_{X}\subseteq U\times\prod_{\alpha\in X}V_{\alpha}$ is such that:
@@
[…]
@@
$=f_{Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in Y{\rangle}{\rangle})=f_{Y}(\pi_{X,Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}))$
@@
[…]
@@ 2. Two properties of filter
Given set $\mathcal{F}\subseteq P(X)$, we denote by *=$\mathcal{F}^{*}=\{X\setminus A\mid A\in\mathcal{F}\}$. When $\mathcal{F}$ is a filter, $\mathcal{F}^{*}$ is an ideal which we call the dual ideal, and when $\mathcal{I}$ is an ideal $\mathcal{I}^{*}$ is a filter which we call the dual filter. Ideals are always considered with the (regular) inclusion order.
@@
[…]
@@ 3. $\underline{(3)\Rightarrow(1)}$
[…]
@@
Let $A_{n}$ be a partition of $\omega$ so that $A_{n}\in U_{n}$. Define $f({\langle}B_{n}\mid n<\omega{\rangle})=\bigcup_{n<\omega}B_{n}\cap A_{n}$. It is clearly monotone. If $X\in\bigcap_{n<\omega}U_{n}$, we let $X\cap A_{n}=B_{n}\in U_{n}$ Then $X=X\cap\omega=X\cap(\bigcup_{n<\omega}A_{n})=\bigcup_{n<\omega}X\cap A_{n}=f({\langle}B_{n}\mid n<\omega{\rangle})$. To see it is unbounded, suppose that $f^{\prime\prime}\mathcal{A}$ is bounded by $B\in\bigcap_{n<\omega}U_{n}$, then for every $A\in\mathcal{A}$, $A\cap A_{n}\supseteq B\cap A_{n}$. and therefore ${\langle}B\cap A_{n}\mid<\omega{\rangle}$ would bound $\mathcal{A}$. ∎
@@ 2.2. Simple and deterministic ideals
$=\bigcup_{a\in\mathcal{A}}X\setminus\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]=\bigcup_{a\in\mathcal{A}}a$
@@
Suppose that $I^{\omega}\equiv_{T}I$, $I$ is deterministic. Then for every ultrafilter $U$ such that $I\subseteq U^{*}$ and $U$ satisfies the $I$-p.i.p, $U\cdot U\equiv_{T}U$.
@@
Suppose that $\mathrm{fin}\subseteq I$ be any ideal over $\omega$, and ${\langle}J_{n}\mid n<\omega{\rangle}$ is a sequence of deterministic ideals over $\omega$ such that for every $n<\omega$, $J_{n+1}\geq_{T}J_{n}$. Then $\sum_{I}J_{n}$ is deterministic.
@@
Let $\mathcal{B}_{n}\subseteq J_{n}$ witness that $J_{n}$ is deterministic. Let ${\langle}f_{m,n}:J_{n}\rightarrow J_{m}\mid n\leq m<\omega{\rangle}$ be a sequence of unbounded maps. Denote by $\mathcal{B}$ the set of all sequences $\vec{b}={\langle}b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n<m<\omega$, $b_{m}\supseteq f_{m,n}(b_{n})$. For $\vec{b}\in\mathcal{B}$ and $A\in I$ defined
@@
Let us show that $\mathcal{C}=\{C_{A,\vec{b}}\mid A\in I,\vec{b}\in\mathcal{B}\}$ is cofinal in $\sum_{I}J_{n}$. Let $Z\in\sum_{I}J_{n}$, and $A=\{n\mid(Z)_{n}\notin J_{n}\}$. By definition of $\sum_{I}J_{n}$, $A\in I$. Construct an increasing sequence $\langle b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n\notin A$, $(Z)_{n}\subseteq b_{n}$ and for every $n<m<\omega$, $f_{m,n}(b_{n})\subseteq b_{m}$. It is possible to construct such a sequence recursively. At stage $n$, we need to find $b_{n}$ such that for all $k<n$, $f_{n,k}(b_{k})\subseteq b_{n}$ and if $n\notin A$, then also $(Z)_{n}\subseteq b_{n}$. These are finitely many sets in $J_{n}$ and therefore, since $\mathcal{B}_{n}$ is a cofinal set in $J_{n}$, we can find a single $b_{n}\in\mathcal{B}_{n}$ including the union of these sets. It follows that for every $n$, $(Z)_{n}\subseteq(C_{A,{\vec{b}}})_{n}$ and therefore $Z\subseteq C_{A,\vec{b}}$. Let us prove that $\mathcal{C}$ witnesses that $\sum_{I}J_{n}$ is deterministic. Suppose that $\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}}\notin\sum_{I}J_{n}$. Then $A:=\{n<\omega\mid(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}\notin J_{n}\}\notin I$. Note that $(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}$. If there is $i_{0}\in T$ such that $n\in X_{i_{0}}$, then $(C_{X_{i_{0}},\vec{b}^{i_{0}}})_{n}=\omega$ and in particular $\omega=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}$. Otherwise, consider $n\in A$ such that for every $i\in T$, $n\notin X_{i}$. Then $(C_{X_{i},\vec{b}^{i}})_{n}\in\mathcal{B}_{n}$ for every $i\in T$ and $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\notin J_{n}$. Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J^{*}_{n}$. Fix any $n_{0}\in A$, which exists  [^17] If $I$ is principle then we pick $n_{0}$ such that $\{n_{0}\}\in I^{+}$. as $A\notin I$. For every $n\geq n_{0}$, either there is $i\in T$ such that $n\in X_{i}$, and as we have seen, $\cup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}=\omega\in J^{*}_{n}$. Otherwise, for every $i\in T$, $(C_{X_{i},\vec{b}^{i}})_{n}=b^{i}_{n}\in\mathcal{B}_{n}$, and by the assumption, $f_{n,n_{0}}(b^{i}_{n_{0}})\subseteq b^{i}_{n}$. Since $n_{0}\in A$, $\bigcup_{i\in T}b^{i}_{n_{0}}\notin J_{n_{0}}$, and since $f_{n,n_{0}}$ is unbounded, $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\notin J_{n}$. Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, it follows that $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\in J^{*}_{n}$. We conclude that for every $n\geq n_{0}$, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J^{*}_{n}$. Since $\{n_{0},n_{0}+1,...\}\in I^{*}$, we have that $\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}}\in(\sum_{I}J_{n})^{*}$ as wanted.
@@
Suppose that $I,J$ are ideals over $\omega$ such that $J$ is deterministic. Then $I\otimes J$ is deterministic.
@@
For every $I$, $I\otimes\mathrm{fin}$ is deterministic.
@@
For any non-principal ultrafilter $U$ such that $U\equiv_{T}U\cdot U$ and for every $I\subseteq U^{*}$, $I^{\omega}\leq_{T}U$.
@@
$I\leq_{T}I\otimes\mathrm{fin}\leq_{T}U\cdot U$. Hence by Corollary 1.8
$I^{\omega}\leq_{T}(U\cdot U)^{\omega}\equiv_{T}(U\cdot U)\cdot(U\cdot U)\equiv_{T}U\cdot U\equiv_{T}U.$
3. Proof of the main result and some corollaries
Let us turn to the proof of the main result which appears in Theorem 3.2 below.

### v2

@@ 1.2. The cofinal type of sums and products
[…]
@@
Suppose that $\mathbb{P}$ is a complete order. Then $\mathbb{P}$ is uniformly below $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\lambda{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
@@
and that $X_{\alpha}\subseteq(B)_{\alpha}$. It follows by monotonicity of the functions, and by $({\dagger})$ that
@@
[…]
@@
Note that if $B^{\prime}\in\mathcal{X}_{X}$ is such that $B^{\prime}\subseteq A$, then the set $\{f_{X}(B)\mid A\subseteq B\in\mathcal{X}_{X}\}$ is bound in $\mathbb{P}$ by $f_{X}(B^{\prime})$ (as $f_{X}$ is monotone). Hence $f^{*}_{X}(A)$ is well defined by completeness of $\mathbb{P}$. It is straightforward that Since $f_{X}$ is monotone cofinal, $f^{*}_{X}$ is monotone cofinal. To see $({\dagger})$, suppose that $Y\subseteq X$, and $C\in U\times\prod_{\alpha<\lambda}V_{\alpha}$, then for every $C\subseteq B\in\mathcal{X}_{X}$, then by $(3)$ $\pi_{X,Y}(C)\subseteq\pi_{X,Y}(B)\in\mathcal{X}_{X}$ and $f_{Y}(\pi_{X,Y}(B))\geq_{\mathbb{P}}f_{X}(B)$. If follows that
@@
$f_{Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in Y{\rangle}{\rangle})=f_{Y}(\pi_{X,Y}({\langle}Z,{\langle}A_{\alpha}\mid\alpha\in X{\rangle}{\rangle}))$
@@
and therefore $f_{X}$ is cofinal. Similarly, to see $({\dagger})$, if $Y\subseteq X\subseteq X_{0}$, and ${\langle}Z,{\langle}A_{n}\mid n\in X{\rangle}{\rangle}\in\mathcal{X}_{X}$, then $\sigma_{Y,X_{0}}^{-1}(n)\geq\sigma_{X,X_{0}}^{-1}(n)$, and therefore, for every $n<\omega$,
@@ 2. Two properties of filter
Given set $\mathcal{F}\subseteq P(X)$, we denote by $\mathcal{F}^{*}=\{X\setminus A\mid A\in\mathcal{F}\}$. When $\mathcal{F}$ is a filter, $\mathcal{F}^{*}$ is an ideal which we call the dual ideal, and when $\mathcal{I}$ is an ideal $\mathcal{I}^{*}$ is a filter which we call the dual filter. Ideals are always considered with the (regular) inclusion order.
@@
    $F$ has $\{\emptyset\}$-p.i.p if and only if $F$ is $\sigma$-complete (and therefore, if $F$ is on $\omega$, then it is principal).
@@ 3. $\underline{(3)\Rightarrow(1)}$
$U_{F}=\{S\in P(F^{*})\mid\exists T\subseteq S\text{ downwards closed }F\text{ has }T\text{-p.i.p}\}.$
@@
Let $A_{n}$ be a partition of $\omega$ so that $A_{n}\in U_{n}$. Define $f({\langle}B_{n}\mid n<\omega{\rangle})=\bigcup_{n<\omega}B_{n}\cap A_{n}$. It is clearly monotone. If $X\in\bigcap_{n<\omega}U_{n}$, we let $X\cap A_{n}=B_{n}\in U_{n}$ Then $X=X\cap\omega=X\cap(\bigcup_{n<\omega}A_{n})=\bigcup_{n<\omega}X\cap A_{n}=f({\langle}B_{n}\mid n<\omega{\rangle})$. To see it is unbounded, suppose that $f^{\prime\prime}\mathcal{A}$ is bounded by $B\in\bigcap_{n<\omega}U_{n}$, then for every $\in\mathcal{A}, AA$$\cap A_{n}\supseteq B\cap A_{n}$. and therefore ${\langle}B\cap A_{n}\mid<\omega{\rangle}$ would bound $\mathcal{A}$. ∎
@@ 2.2. Simple and deterministic ideals
$\bigcup_{a\in\mathcal{A}}X\setminus\pi^{-1}[\pi^{\prime\prime}[X\setminus a]]=\bigcup_{a\in\mathcal{A}}a$
@@
Suppose that $^{\omega}\equiv_{T}I, II$ is deterministic. Then for every ultrafilter $U$ such that $I\subseteq U^{*}$ and $U$ satisfies the $I$-p.i.p, $U\cdot U\equiv_{T}U$.
@@
Suppose that $\mathrm{fin}\subseteq I$ be a deterministic ideal over $\omega$, and ${\langle}J_{n}\mid n<\omega{\rangle}$ is a sequence of deterministic ideals over $\omega$ such that for every $n<\omega$, $J_{n+1}\geq_{T}J_{n}$. Then $\sum_{I}J_{n}$ is deterministic.
@@
Let $\mathcal{B}_{n}\subseteq J_{n}$ witness that $J_{n}$ is deterministic, and $\mathcal{D}$ witness that $I$ is deterministic. Let ${\langle}f_{m,n}:J_{n}\rightarrow J_{m}\mid n\leq m<\omega{\rangle}$ be a sequence of unbounded maps. Denote by $\mathcal{B}$ the set of all sequences $\vec{b}={\langle}b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n<m<\omega$, $b_{m}\supseteq f_{m,n}(b_{n})$. For $\vec{b}\in\mathcal{B}$ and $A\in\mathcal{D}$ defined
@@
Let us show that $\mathcal{C}=\{C_{A,\vec{b}}\mid A\in\mathcal{D},\vec{b}\in\mathcal{B}\}$ is cofinal in $\sum_{I}J_{n}$. Let $Z\in\sum_{I}J_{n}$, and $A=\{n\mid(Z)_{n}\notin J_{n}\}$. By definition of $\sum_{I}J_{n}$, $A\in I$, so there is $D\in\mathcal{D}$ such that $A\subseteq D$. Construct an increasing sequence $\langle b_{n}\mid n<\omega{\rangle}\in\prod_{n<\omega}\mathcal{B}_{n}$ such that for every $n\notin A$, $(Z)_{n}\subseteq b_{n}$ and for every $n<m<\omega$, $f_{m,n}(b_{n})\subseteq b_{m}$. It is possible to construct such a sequence recursively. At stage $n$, we need to find $b_{n}$ such that for all $k<n$, $f_{n,k}(b_{k})\subseteq b_{n}$ and if $n\notin A$, then also $(Z)_{n}\subseteq b_{n}$. These are finitely many sets in $J_{n}$ and therefore, since $\mathcal{B}_{n}$ is a cofinal set in $J_{n}$, we can find a single $b_{n}\in\mathcal{B}_{n}$ including the union of these sets. It follows that for every $n$, $(Z)_{n}\subseteq(C_{D,{\vec{b}}})_{n}$ and therefore $Z\subseteq C_{D,\vec{b}}$. Let us prove that $\mathcal{C}$ witnesses that $\sum_{I}J_{n}$ is deterministic. Suppose that $\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}}\notin\sum_{I}J_{n}$. Then $A:=\{n<\omega\mid(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}\notin J_{n}\}\notin I$. Note that $(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}$. Let us split into cases: If $A\subseteq\bigcup_{i\in T}X_{i}$, then $\bigcup_{i\in T}X_{i}\notin I$ and since $\{X_{i}\mid i\in T\}\subseteq\mathcal{D}$, $\bigcup_{i\in T}X_{i}\in I^{*}$. Now for every $n\in\bigcup_{i\in T}X_{i}$, there is $i_{0}\in T$ such that $(C_{X_{i_{0}},\vec{b}^{i_{0}}})_{n}=\omega$ and in particular $\omega=\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}$. We conclude that $\{n<\omega\mid(\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}})_{n}\in J_{n}^{*}\}\supseteq\bigcup_{i\in T}X_{i}\in I^{*}$. Hence $\bigcup_{i\in T}C_{i}\in(\sum_{I}J_{n})^{*}$.
Otherwise, consider $n_{0}\in A$ such that for every $i\in T$, $n_{0}\notin X_{i}$. Then $(C_{X_{i},\vec{b}^{i}})_{n_{0}}\in\mathcal{B}_{n_{0}}$ for every $i\in T$ and $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n_{0}}\notin J_{n_{0}}$. Since $\mathcal{B}_{n_{0}}$ witnesses that $J_{n_{0}}$ is deterministic, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n_{0}}\in J^{*}_{n_{0}}$. For every $n\geq n_{0}$, either there is $i\in T$ such that $n\in X_{i}$, and as we have seen, $\cup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}=\omega\in J^{*}_{n}$. Otherwise, for every $i\in T$, $(C_{X_{i},\vec{b}^{i}})_{n}=b^{i}_{n}\in\mathcal{B}_{n}$, and by the assumption, $f_{n,n_{0}}(b^{i}_{n_{0}})\subseteq b^{i}_{n}$. Since $\bigcup_{i\in T}b^{i}_{n_{0}}\notin J_{n_{0}}$, and $f_{n,n_{0}}$ is unbounded, $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\notin J_{n}$. Since $\mathcal{B}_{n}$ witnesses that $J_{n}$ is deterministic, it follows that $\bigcup_{i\in T}f_{n,n_{0}}((C_{X_{i},\vec{b}^{i}})_{n})\in J^{*}_{n}$. We conclude that for every $n\geq n_{0}$, $\bigcup_{i\in T}(C_{X_{i},\vec{b}^{i}})_{n}\in J^{*}_{n}$. Since $\{n_{0},n_{0}+1,...\}\in I^{*}$, we have that $\bigcup_{i\in T}C_{X_{i},\vec{b}^{i}}\in(\sum_{I}J_{n})^{*}$ as wanted.
@@
Suppose that $I,J$ are deterministic ideals over $\omega$. Then $I\otimes J$ is deterministic.
Recall that for an ideal $I$ on $\omega$ we define recursively the ideal $I^{\otimes\alpha}$ on $\omega^{\alpha}$ by setting $I^{1}=I$, $I^{\otimes(\alpha+1)}=I\otimes I^{\otimes\alpha}$ on $\omega^{\alpha+1}=\omega\times\omega^{\alpha}$. For limit $\alpha<\omega_{1}$, fix a cofinal sequence ${\langle}\alpha_{n}\mid n<\omega{\rangle}$ in $\alpha$, then $I^{\otimes\alpha}=\sum_{I}I^{\otimes\alpha_{n}}$ is an ideal on $\bigcup_{n<\omega}\{n\}\times\omega^{\alpha_{n}}$. It is clear that $\alpha\leq\beta$ then $I_{\alpha}\leq_{RK}I_{\beta}$ (see for example [3, Lemma 3.2]).
@@
If $I$ is deterministic then for every $\alpha<\omega_{1}$, $I^{\otimes\alpha}$ is deterministic. In particular $\mathrm{fin}^{\otimes\alpha}$ is deterministic for every every $\alpha<\omega_{1}$.
Proof.
By induction on $\alpha$. For $\alpha=1$ II is deterministic by assumption. Suppose that this was true for $\alpha$, then $I^{\alpha+1}=I\otimes I^{\otimes\alpha}$ which is deterministic from the previous corollary. Similarily, at limit steps we just use the induction hypothesis and Proposition 2.29.
In [3] it was noticed that a generic ultrafilter for $P(X)/I$ where $I$ is a $\sigma$-ideal satisfies the $I$-p.i.p. now together with Corollary 2.27, we recover the result from [3] in our abstract settings.
Fact 2.32.
For every $\alpha\geq 2$, $I^{\otimes\alpha}\equiv_{T}I\otimes I\equiv_{T}I^{\omega}$.
Proof.
By induction on $\alpha\geq 2$. For $\alpha=2$,
$(I\otimes I)^{\omega}\equiv_{T}(I^{\omega})^{\omega}\simeq I^{\omega}\equiv_{T}I\otimes I.$
For successor $\alpha+1$, we use that fact that $I\otimes I\leq_{T}I^{\otimes\alpha}$ to see that
$I^{\otimes(\alpha+1)}\equiv_{T}I\otimes I^{\otimes\alpha}\equiv_{T}I\otimes(I\otimes I)\equiv_{T}I\otimes I$
At limit steps we have that
$I^{\otimes\alpha}\equiv_{T}\sum_{I}I^{\otimes\alpha}\equiv_{T}\sum_{I}I\times I\equiv_{T}I\otimes(I\otimes I)\equiv_{T}I\otimes I$
In particular, $(I^{\otimes\alpha})^{\omega}\equiv_{T}I^{\otimes\alpha}$.
@@
Let $I$ be a deterministic $\sigma$-ideal. Then for every $2\leq\alpha<\omega_{1}$, a generic ultrafilter $G$ for $P(\omega^{\alpha})/I^{\otimes\alpha}$ satisfies $G\cdot G\equiv_{T}G$.
@@
$G$ has the $I^{\otimes\alpha}$-p.i.p, and by Corollary 2.31 $I^{\otimes\alpha}$ is deterministic. Also by the previous fact, $(I^{\otimes\alpha})^{\omega}\equiv_{T}I^{\otimes\alpha}$ . Thus by Corollary 2.27 $G\cdot G\equiv_{T}G$. ∎
For $\alpha=1$ we can also say something:
Proposition 2.34.
Let $I$ be a simple $\sigma$-ideal. If $G\subseteq P(\omega)/I$ is $V$-generic then $G\cdot G\equiv_{T}G\times I^{\omega}$. So in particular if $I^{\omega}\equiv_{T}I$, $G\cdot G\equiv_{T}G$
Proof.
Since $G$ has the $I$-p.i.p, $G\cdot G\leq_{T}G\times I^{\omega}$. In the other direction, we known that $I^{*}\subseteq G$ and since $I$ is simple, $I\leq_{T}G$. Hence $I^{\omega}\leq_{T}G^{\omega}\equiv_{T}G\cdot G$. We conclude that $G\times I^{\omega}\leq G\cdot G$. ∎
3. Commutativity of cofinal types
Definition 3.1.
We say that a class $\mathcal{C}$ of ultrafilters on $\omega$ is a commutative class if for every $U,W\in\mathcal{C}$, $U\cdot W\equiv_{T}W\cdot U$.
Example 3.2.
By Dobrinen and Todorcevic, the class $\mathcal{C}_{RP}$ of rapid $p$-points is a commutative class and if $W$ is a rapid $p$-point then $U\cdot W\equiv_{T}U\times W$.
Example 3.3.
By Milovich [29], the class $\mathcal{C}_{P}$ of $p$-points is a commutative class since if $W$ is a $p$-point, then $U\cdot W\equiv_{T}U\times W\times\omega^{\omega}$.
Fact 3.4.
For any three ultrafilters $U,W,Z$, $U\cdot(W\cdot Z)\equiv_{RK}(U\cdot W)\cdot Z$
From the above associativity, we get the following:
Proposition 3.5.
If $\mathcal{C}$ is a commutative class then so is $\overline{\mathcal{C}}=\{U_{1}\cdot U_{2}...\cdot U_{n}\mid U_{1},...,U_{n}\in\mathcal{C}\}$.
One use of commutative classes relates to the character of an ultrafilter. Recall that for an ultrafilter $U$, $Ch(U)$ denoted the minimal size of a cofinal subset of $U$. It is clear that if $U\equiv_{T}U^{\prime}$ then $Ch(U)=Ch(U^{\prime})$. Hence we conclude the following: